OpenStudy (anonymous):

Simplify the following expression: a + 2c - 5b - c + a - b a.2a - 6b + 2c b.2a - 6b + c c.a - 6b + c d.a - 5b + 2c

OpenStudy (anonymous):

we have a+a = 2a and 2c-c = c and -5b-5 = -6b

OpenStudy (anonymous):

got it?

OpenStudy (uri):

a + 2c - 5b - c + a - b Combine all a values,Combine all c values and combine all b values and then add/subract.

OpenStudy (anonymous):

so the answer is: 2a + c -6b

OpenStudy (anonymous):

are you sure @PFEH.1999

OpenStudy (uri):

@jasmin16 Try doing it yourself..He can be wrong...

OpenStudy (anonymous):

yes i'm quite sure about this,but @uri is right...pay attention to my first command!

OpenStudy (uri):

@jasmin16 If you want JUST answers you can go to Wolfram or Weegy.com...Yahoo is best though in giving ANSURS.OpenStudy is to help.Please don't break the rules.

OpenStudy (anonymous):

I think it is a.

OpenStudy (uri):

Well solve it and show your work?

OpenStudy (uri):

We told you how to do it.

OpenStudy (anonymous):



OpenStudy (anonymous):

I think it's a good way to teach :)